OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –4? A. y = 3x + 4 B. y = –3x – 4 C. y = –3x + 4 D. y = 3x – 4

OpenStudy (anonymous):

@imawesome101

OpenStudy (anonymous):

In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So what do you think the answer is ?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

look again....you got the slope right, but look at your y intercept again

OpenStudy (anonymous):

I would have to say c

OpenStudy (anonymous):

b then thats what i say and the what the fox says lol :P

OpenStudy (anonymous):

let me show you slope(m) = -3 y intercept (b) = -4 y = mx + b now we sub y = -3x + (-4) y = -3x - 4 <-- answer B

OpenStudy (anonymous):

So it would have to be B right

OpenStudy (anonymous):

yes..it is B

OpenStudy (anonymous):

thats what the fox says...lol...good one

OpenStudy (anonymous):

thanks and bobbie either ur sayin boobie or bobie wich one

OpenStudy (anonymous):

i know when i said thats what i say i thought of what does the fox say

OpenStudy (anonymous):

I think it is Bobby with an ie instead of a y

OpenStudy (anonymous):

o

OpenStudy (anonymous):

:O :) :( :P :D :C